OpenStudy (anonymous):

find the parametric equation of the following curve: ellipse x^2+y^2=1 , z=y

OpenStudy (loser66):

I have question: is x^2 + y^2 =1 an ellipse? is it not a circle?

OpenStudy (anonymous):

no its written as ellipse in the qst

OpenStudy (anonymous):

x^2 + y^2 =1 is a circle. Look: https://www.wolframalpha.com/input/?i=x%C2%B2%2By%C2%B2%3D1&lk=4&num=2

OpenStudy (loser66):

@IgorL I think so, too. but you see, he answered that is an ellipse. Whatever, we have to find out the parametric equation for that, (anyway, circle is a special case of ellipse) . I stuck at z = y and try to find out the parametric for that part.

OpenStudy (experimentx):

x = y => x(t) = t, y(t) = t

OpenStudy (anonymous):

ok

OpenStudy (loser66):

@experimentX no, z = y not x =y

OpenStudy (loser66):

x = cos t and y = sin t

OpenStudy (experimentx):

if you are considering 3d space, x^2+y^2=1 , z=y is not a curve anymore, it's a region

OpenStudy (loser66):

yes, I need the pic, too. please, draw it out

OpenStudy (loser66):

@experimentX can we solve it by algebraic way only?

OpenStudy (experimentx):

woops!! it's a curve ... sorry

OpenStudy (experimentx):

i misinterpreted, it's an ellipse indeed

OpenStudy (loser66):

:) don't get

OpenStudy (experimentx):

x^2+y^2 = 1 is a cylinder and x=z is a plane, when a cylinder intersects plane what do you get?

OpenStudy (loser66):

nope, sorry, ellipse only

OpenStudy (experimentx):

no ... it's an ellipse.